OpenStudy (anonymous):

What is the standard form equation of the line passing through (–2, 8) with a slope of 2

OpenStudy (anonymous):

0put into point slope form which is y-y1=M(x-x1) where m is the slope and (x1,y1) represents the given point (-2,8)

OpenStudy (anonymous):

plug it in then put into this form ay+bx=c

OpenStudy (anonymous):

I get everything except the plugging in part

OpenStudy (anonymous):

2x – y = 12 2x + y = 12 Im stuk between these choices i get the point slope formula i just don't know how to write the equation in standard form

hartnn (hartnn):

slope m = 2  got that much ?